OpenStudy (anonymous):

Find the value of y so that so that the line passing through (4, y) and (5, 3) has a slope of 1

OpenStudy (anonymous):

\[\frac {y - 3}{4 - 5} = 1 => y - 3 = -1 => y = 2\]